OpenStudy (anonymous):

solve: 3x/2+1/x=3x/4

OpenStudy (bahrom7893):

= (3x^2 + 2) / (2x) = 3x/4

OpenStudy (bahrom7893):

Cross multiply: \[6x^2 = 4*(3x^2+2)\]

OpenStudy (bahrom7893):

\[6x^2 = 12x^2 + 8\]

OpenStudy (anonymous):

and then subtract the 6x^2 then what?

OpenStudy (bahrom7893):

sorry smth is wrong..checkin my work

OpenStudy (anonymous):

well i got the same thing

OpenStudy (bahrom7893):

then either the question is wrong or no solutions exist..

OpenStudy (bahrom7893):

bcuz u'll get -6x^2 = 8... x^2 cant be negative

OpenStudy (anonymous):

yeah i know..theres no other way of solving it?

OpenStudy (bahrom7893):

NO.. unless u use imaginary number.. which i doubt

OpenStudy (anonymous):

yeaa we doo

OpenStudy (bahrom7893):

-6x^2 = 8 x^2 = -8/6

OpenStudy (anonymous):

alright thank uu go to sleep!

OpenStudy (bahrom7893):

x= i * sqrt(8/6), where i is sqrt(-1)